OpenStudy (anonymous):

find 3 consecutive odd intergers with a sum of 63. show work.

OpenStudy (zarkon):

Solve \[n+(n+2)+(n+4)=63\] for n

OpenStudy (radar):

Let 2x+1 = first odd integer then 2x+3 = next and 2x + 5 = last one 6x+9 = 63 6x=54 x=9 2x+1=19 2x+3=21 2x+5 =23 The answer is 19, 21, and 23

OpenStudy (zarkon):

\[n+(n+2)+(n+4)=63\] \[3n+6=63\] \[3n=57\] \[n=19\] so 19,21,23